You need to prepare 100.0 mL of a pH=4.00 buffer solution using 0.100 M benzoic acid (pKa = 4.20) and 0.140 M sodium benzoate. How much of each solution should be mixed to prepare this buffer?

mL benzoic acid = ?
mL sodium benzoate = ?
Answer #1
Concepts and reason

Concentration of a weak acid is related to its pKa{\rm{pKa}}and pH of a solution as shown in the following Henderson -Hasselbalch equation. The equation is shown below.

pH=pKa+log[salt][acid]{\rm{pH = pKa + log}}\frac{{\left[ {{\rm{salt}}} \right]}}{{\left[ {{\rm{acid}}} \right]}}

Here, Ka is the acid dissociation constant; [salt]\left[ {{\rm{salt}}} \right], and[acid]\left[ {{\rm{acid}}} \right]are the concentration terms of the salt and the acid respectively. This equation is applicable to a system that contains weak acids or weak bases only.

Benzoic acid is a weak acid, so we can apply the equation to this system. Calculate the volumes of benzoic acid and sodium benzoate by substituting the values ofpKa{\rm{pKa}} and pH in the equation.

For the benzoic acid and for the buffer solution, the equation becomes as shown below.

pH=pKa+log[Sodiumbenzoate][Benzoicacid]{\rm{pH = pKa + log}}\frac{{\left[ {{\rm{Sodium benzoate}}} \right]}}{{\left[ {{\rm{Benzoic acid}}} \right]}}

Fundamentals

An aqueous solution that contains a mixture of a weak acid and its conjugate base or vice versa is called as buffer solution.

Benzoic acid is a weak acid. Because of its poor dissociation, it forms benzoate ions. The Henderson -Hasselbalch equation is applicable to weak acids and bases only.

The value of pKa{\rm{pKa}}for the benzoic acid is 4.20. Calculate the fraction of concentrations of acetate and acetic acid.

pH=pKa+log[Sodiumbenzoate][Benzoicacid]4.0=4.20+log[Sodiumbenzoate][Benzoicacid]\begin{array}{c}\\{\rm{pH = pKa + log}}\frac{{\left[ {{\rm{Sodium benzoate}}} \right]}}{{\left[ {{\rm{Benzoic acid}}} \right]}}\\\\4.0 = 4.20 + {\rm{log}}\frac{{\left[ {{\rm{Sodium benzoate}}} \right]}}{{\left[ {{\rm{Benzoic acid}}} \right]}}\\\end{array}

log[Sodiumbenzoate][Benzoicacid]=0.20[Sodiumbenzoate][Benzoicacid]=100.20=0.631\begin{array}{c}\\{\rm{log}}\frac{{\left[ {{\rm{Sodium benzoate}}} \right]}}{{\left[ {{\rm{Benzoic acid}}} \right]}} = - 0.20\\\\\frac{{\left[ {{\rm{Sodium benzoate}}} \right]}}{{\left[ {{\rm{Benzoic acid}}} \right]}} = {10^{ - 0.20}} = 0.631\\\end{array}

Assume the volume of the acid is xx and volume of the benzoate solution asyy.

x+y=100mLx + y = 100{\rm{ mL}}(1)\left( 1 \right)

[Sodiumbenzoate][Benzoicacid]=0.631\frac{{\left[ {{\rm{Sodium benzoate}}} \right]}}{{\left[ {{\rm{Benzoic acid}}} \right]}} = 0.631

y×0.140Mx×0.1M=0.631\frac{{y \times 0.140{\rm{ M}}}}{{x \times 0.1{\rm{ M}}}} = 0.631

y=0.451×xy = 0.451 \times x …. (2)\left( 2 \right)

By solving the equations (1)\left( 1 \right) and(2)\left( 2 \right),

x=68.93mLy=31.07mL\begin{array}{l}\\x = 68.93{\rm{ mL}}\\\\{\rm{y = 31}}{\rm{.07 mL}}\\\end{array}

Ans:

Volume of benzoic acid is68.93mL68.93{\rm{ mL}}.

Volume of sodium benzoate is31.07mL{\rm{31}}{\rm{.07 mL}}.